# cMPI: Using CXL Memory Sharing for MPI One-Sided and Two-Sided Inter-Node Communications

### v1

@@ Abstract.
Message Passing Interface (MPI) is a foundational programming model for high-performance computing. MPI libraries traditionally employ network interconnects (e.g., Ethernet and InfiniBand) and network protocols (e.g., TCP and RoCE) with complex software stacks for cross-node communication. This paper presents cMPI, the first work to optimize MPI point-to-point communication (both one-sided and two-sided) using CXL memory sharing on a real CXL platform, transforming cross-node communication into memory transactions and data copies within CXL memory, bypassing traditional network protocols. We analyze performance across various interconnects and find that CXL memory sharing achieves 7.2-8.1 lower latency than TCP-based interconnects deployed in small- and medium-scale clusters. We address challenges of CXL memory sharing for MPI communication, including data object management over the dax representation (Ma et al., 2024), cache coherence, and atomic operations. Overall, cMPI outperforms TCP over standard Ethernet NIC and high-end SmartNIC by up to 49 and 72 in latency and bandwidth, respectively, for small messages.
@@ 1. Introduction
Given two MPI processes distributed on two nodes, two types of MPI communication can be employed, either one-sided or two-sided depending on process participation. The MPI communication traditionally employs high-performance interconnect (e.g., Ethernet), and relies on Remote Direct Memory Access (RDMA) or network protocols (e.g., RoCE (Schelten et al., 2022)). The support of high-performance interconnect includes a complicated software stack, including application-level protocol implementation and system-level driver for device registration and interrupt handling. Using the interconnect technologies also requires the MPI library to use special APIs or programming methods (e.g., socket programming).
@@
Recent work (Peng et al., 2022; Wahlgren et al., 2023a; Li et al., 2023b) reveals low memory utilization in individual nodes in data centers or (supercomputers), and demonstrates the potential of memory disaggregation for improving memory utilization and reducing production cost on supercomputers (Fridman et al., 2023; Wahlgren et al., 2023a; Li et al., 2023a). The emerging Compute Express Link (CXL) (Sharma et al., 2023) makes memory disaggregation highly feasible. CXL is an interconnect technology designed for high-speed communication among processors, accelerators, and memory devices. CXL allows for memory expansion beyond traditional CPU memory channels by enabling a CPU to communicate with DRAM managed by its own controller over a PCIe channel, hence offering increased capacity and bandwidth without increasing the number of primary CPU memory channels. The CXL memory can be accessed using traditional CPU instructions (e.g., load and store), allowing the existing applications to use it with little changes. The CXL memory has been explored as a memory tier to build HPC systems with larger memory capacity for checkpointing or accommodating scientific applications with larger memory footprint (Wahlgren et al., 2022; Fridman et al., 2023; Giles and Varman, 2024; Wang et al., 2025; Wahlgren et al., 2023a).
@@
In this paper, we study a new MPI communication scheme using CXL memory sharing. The CXL memory sharing is based upon the CXL pooled memory platform (a solution for memory disaggregation), shown in Figure 1. Using the CXL memory sharing, the MPI communication across nodes can be transformed into regular memory transactions and data copy within the CXL memory, without going through the traditional network protocols. There are a couple of benefits of doing so. First, the communication latency can be shorter, compared with network‑based approaches such as TCP over Ethernet and RoCEv2. For example, given a message size of 8 bytes, we show that using a real CXL pooled-memory platform (not based on hardware simulation or emulation (Ahn et al., 2022; Wu et al., 2024; Laboratory, [n. d.])) to access the CXL memory takes only 790 , while using TCP over a standard Ethernet NIC and TCP over Mellanox (CX-6 Dx) to access data across nodes take 16 and 18 respectively, about an order of magnitude longer than using the CXL memory sharing (see Section 2 for details). Second, using the CXL memory sharing leads to a simpler software stack, compared to using the traditional network protocols. This is because the CXL protocol to perform CXL memory transactions for data (de)packaging and synchronization between CPU and memory device are done by hardware (not system software).
@@
However, using the CXL memory sharing for MPI communication across nodes faces challenges. First, the representation of the CXL memory sharing in the host (node) creates challenges to manage data objects in the CXL memory. At the node, the CXL memory sharing is exposed as a Direct Access (dax) device by the CXL driver and daxctl (Ma et al., 2024). The dax device bypasses the traditional page cache and provides memory-level access. However, the abstract of the dax device cannot provide flexibility to create arbitrary data objects and manage their life cycle (e.g., creation and deletion).
@@
We develop a library, cMPI, by extending MPICH-4.2.3 (Lab, [n. d.]). We address the limitation of CXL memory sharing to support MPI communications.
@@ 2.2. CXL
The recent CXL specification (CXL 3.0) introduces CXL memory sharing. Based upon the Multi-Headed Device (MHD) and Dynamic Capacity Device (DCD), a host (node) can connect to the CXL pooled memory platform through a dedicated CXL port, ensuring bandwidth fairness and eliminating added latency of a switch hop, shown in Figure 1. We use a CXL pooled memory platform, Niagara 2.0 (Choi, [n. d.]), for CXL memory sharing.
@@
To measure latency and bandwidth, we use Intel MLC (Intel Corporation, 2019) for the cases (1) and (7), iPerf (Laboratory, 2025) for (2) and (3), and Perftest (Distribution, [n. d.]) for (4). We choose those different tools (benchmarks) based on whether they are aligned with the specific system features (e.g., RDMA, or TCP over Ethernet). As we do not have Mellanox CX-3 and CX-6, the performances for the cases (5) and (6) are taken from the vendor’s product report (Technologies, [n. d.]; NVIDIA, [n. d.]). For the case (8), since existing tools do not support testing memory access with cache flushing on a dax device (the system-level abstraction for CXL memory sharing), we develop a micro-benchmark. This benchmark uses mmap to obtain the virtual address space of the dax device. For latency tests, this benchmark performs memset with cache flushing. During the test, we vary the data size and measure the average latency over 1000 iterations. For bandwidth tests, this benchmark uses multiple threads, and each thread accesses 512 bytes with cache flushing on a dax device region. We measures aggregated bandwidth. See Table 1 for results.
@@ 3.1. CXL SHM Arena
Multi-level hashing. For efficient SHM object indexing, we employ a hash-based structure rather than a plain array to construct the meta-data region. We design the hash structure based on two requirements. First, it should effectively resolve hash collisions while maintaining a fixed capacity to avoid dynamic resizing. Second, it must support parallel insertions and searches. To meet these requirements, we use an existing fixed-capacity multi-level hashing scheme (Broder and Karlin, 1990), which is a representative of many multi-level hashing techniques (Broder and Karlin, 1990; Broder and Mitzenmacher, 2001; Zuo et al., 2019; Xiong et al., 2023; Zuo et al., 2018; Chen et al., 2020).
@@
Famfs vs. CXL SHM Arena. Famfs (CXL Micron Research Kit, 2024) is a shared-memory file system framework recently created by MICRON. Famfs is different from our design from three perspectives. First, Famfs lives in the Linux kernel. In contrast, CXL SHM Arena is a user-space library without any kernel modification. Second, Famfs uses a client/master architecture, and restricts the creation of shared files exclusively to the master node, limiting its applicability to MPI where any node may need to create SHM objects. In contrast, CXL SHM Arena supports the creation of SHM objects by arbitrary nodes. Third, the SHM management APIs in CXL SHM Arena resemble those in the traditional POSIX SHM, which facilitates easy integration into the application, whereas Famfs uses quite different APIs.
@@ 3.5. Memory Coherence
CXL SHM allows multiple hosts to access the same memory region. However, our CXL SHM lacks the cache coherence support, which causes the updates made by one host to be invisible to others. To maintain data coherence, we must rely on either non-cacheable memory sharing or a software-based cache coherence mechanism. While CXL 3.0 introduces support for Back-Invalidate (BI) to enforce hardware coherence across hosts, the anticipated growth in CXL memory size makes impractical a precise snoop filter that tracks each cache line (Jain et al., 2024). In contrast, the software-based cache coherence has been shown to succeed to optimize performance in large shared address space (Sun et al., 2023; Ma et al., 2024). We discuss non-cacheable memory sharing and software-based cache coherence as follows.
@@
Software-based cache coherence. We use cache flush and memory fence. Cache flush instructions, such as clwb, clflush, and clflushopt, invalidate the cache line for specific addresses from each level of the cache hierarchy in the coherence domain; if that cache line contains modified data, that data is written back to memory (Guide, 2011). Memory fence instructions, such as store fence (sfence) and load fence (lfence), enforce ordering constraints on memory operations, ensuring that all operations issued prior to the fence are completed before those issued after the fence. We perform cache flush and memory fence when accessing the CXL SHM Arena, CXL SHM-based RMA windows, and CXL SHM-based message queues. Specifically, after every write, we perform a cache flush followed by a memory fence to guarantee that data is written back to memory. Conversely, before every read, we perform a memory fence followed by a cache flush to ensure that stale data, potentially held in caches or due to prefetching, is invalidated.
@@ 3.6. Discussion
cMPI currently focuses on one-sided and two-sided communications across nodes. There are collective communications such as Allgather, Allreduce, and ReduceScatter that can benefit from CXL SHM. Within an MPI library, these collective communications are often implemented using point-to-point communications based on certain algorithms (e.g., recursive doubling (Arap et al., 2015) and Bruck’s algorithm (Fan et al., 2022)), hence the collective communications can directly benefit from cMPI. Nevertheless, there are multiple challenges of using cMPI for the collective communicators, such as communicator handling and integrating intra/inter-node point-to-point communications. We leave collective communications as future work.
@@ 3.7. More Implementation Details
We implement cMPI by extending MPICH-4.2.3 (Gropp, 2002). In CXL SHM Arena, to enable cMPI to create millions of distinct CXL SHM objects, we configure a 10-level multi-level hash table and set the maximum number of slots at the level one to 200,000. Since the number of slots is constrained to be a prime number, the actual slot counts across the levels 1-10 range from 199,999 to 199,873. This results in a total of 1,999,260 slots across all levels.
@@ 4.1. Evaluation Methodology
Real CXL Platform. We use Niagara 2.0 (Choi, [n. d.]), an FPGA-based CXL pooled memory platform. This platform can connect up to 4 nodes without a CXL switch. Each server connects to this platform via an MCIO x8 cable using PCIe 4.0 (16 GT/s per lane). The platform uses four DDR4-2400 DIMM channels, with a total capacity of 1TB and a bandwidth of 19.2 GB/s per channel.
@@
Simulator for scalability study. We use SimGrid (Casanova et al., 2025), an event-based simulator that can study MPI applications with user-specified interconnect latency and bandwidth. Because our CXL platform is constrained to connecting only four hosts, simulation is necessary to study scalability beyond this hardware limitation.
@@ 4.2. Microbenchmarks
We use the OSU Micro-Benchmark suite (OMB) (Laboratory, [n. d.]) to compare performance (i.e., latency and bandwidth) across two nodes using CXL SHM, TCP over Ethernet, and TCP over Mellanox (CX-6 Dx). To evaluate the performance of one-sided communication under multiple processes, we extend the OMB’s benchmark for one-sided communication, which originally only supports two processes, to support any number of MPI origin and target processes. To evaluate the performance of two-sided communication, we change the size per message cell to 64KB to improve performance (see Section 4.3). We evaluate the message size from 1B to 8MB, representing small, medium, and large message sizes (Laboratory, [n. d.]). For one-sided communication, given MPI processes, half of them is used as the origin processes, and the other half is used as the target processes.
@@ 4.4. Scalability Study
We use CG (a conjugate gradient solver) from NAS Parallel Benchmark (NPB) Suite (Bailey et al., 1991) 3.4. For CG, we use Class D as input. We also use miniAMR (Vaughan and Barrett, 2015) (a proxy application for adaptive mesh refinement) with the input block size 4 in directions. We choose CG and miniAMR for evaluation, because their communications are dominated by MPI two-sided communications (Sultana et al., [n. d.]; Small and Yuan, 2009).
@@ 5. Related Work
Intra-node MPI communication over shared memory. Optimizing intra-node MPI communication within shared memory has been extensively studied (Li et al., 2018; Graham and Shipman, 2008; Huang et al., 2008; Adam et al., 2025; Walia et al., 2024; Cho et al., 2017; White and Kale, 2020; Huang et al., 2003). For example, Adam et al. (Adam et al., 2025) reveal significant bandwidth differences between cores and sockets in a node and demonstrate that current MPI implementations may not fully utilize the intra-node bandwidth potential. Walia et al. (Walia et al., 2024) leverage OpenMP for accelerating MPI_Allreduce in deep learning workloads. Cho et al. (Cho et al., 2017) improve MPI bandwidth performance through on-package memory and huge page support. White and Kale (White and Kale, 2020) optimize point-to-point communication in AMPI (Huang et al., 2003) by their locality-aware design in shared memory. Unlike previous works focusing on intra-node communication, our study optimizes inter-node MPI communication over CXL SHM.
@@
Evaluation of CXL. Recent CXL memory evaluations provide insights into this emerging memory technology (Liu et al., 2024; Sun et al., 2023; Yang et al., 2025a; Wang et al., 2025; Sun et al., 2023; Tang et al., 2024; Wu et al., 2024). For example, Sun et al. (Sun et al., 2023) demonstrate important differences between real and emulated CXL memory, necessitating reconsideration of previous assumptions on the CXL hardware. Tang et al. (Tang et al., 2024) evaluate CXL performance and design an Abstract Cost Model to estimate the cost-benefit of using CXL memory in datacenters. Wang et al. (Wang et al., 2025) provide a comprehensive evaluation of how CXL interacts with GPU in LLM training and inference.
@@
CXL shared memory management. Famfs (CXL Micron Research Kit, 2024; Aravind and John, 2024) from Micron is an open source file system for scale-out disaggregated shared memory across multiple nodes. We discuss its difference from cMPI in Section 3.1. Zhang et al. (Zhang et al., 2023) develop a failure-resilient CXL SHM system using reference counting to tolerate partial client failures.
@@
CXL for MPI. OMB-CXL (Tran et al., 2024) introduces a micro-benchmark suite for evaluating MPI communications over CXL SHM based on emulation. Ahn et al. (Ahn et al., 2024) present the first MPI over CXL SHM implementation beyond benchmarking, utilizing CXL SHM to accelerate MPI_Allgather across nodes. However, this work addresses only a single collective operation and is based on emulation. In contrast, our work is the first to optimize MPI point-to-point communication (both one-sided and two-sided) on real CXL Pooled Memory Platforms, with novel approaches to build DAX abstraction, MPI synchronization, and memory coherence.
@@
Applications of CXL memory. CXL has been explored by various application domains. For databases, Ahn et al. (Ahn et al., 2022) leverage CXL memory expansion to enhance in-memory database systems. For AI workloads, Xie et al. (Xie et al., 2024) introduce SmartQuant, a CXL-based AI model store supporting runtime weight quantization to reduce inference latency. For approximate nearest neighbor search (ANNS), Jang et al. (Jang et al., 2023) propose CXL-ANNS, which utilizes CXL-based memory disaggregation and parallel search techniques to efficiently handle large datasets.
@@
Tiered memory systems. Apart from serving as a shared memory pool, CXL can also be used to build a tiered memory system for a host. With a tiered memory system, various memory components (Wang et al., 2020; Xiang et al., 2022; Bergman et al., 2022; Dragojević et al., 2014; Gu et al., 2017; Abulila et al., 2019; Van Doren, 2019; Sharma, 2022; Chen et al., 2014; Wang et al., 2025) exhibit differences in latency, bandwidth, capacity, and cost. A range of solutions (Maruf et al., 2022; Kumar et al., 2021; Hildebrand et al., 2020; Li and Wu, 2022; Ren et al., 2021b; Wang et al., 2019; Dulloor et al., 2016; Raybuck et al., 2021; Kannan et al., 2017; Kim et al., 2021; Weiner et al., 2022; Wu et al., 2018, 2017; Ren et al., 2021a; Wahlgren et al., 2023b; Xie et al., 2023; Ren et al., 2024; Xu et al., 2024; Liu et al., 2018; Huang and Li, 2017; Ren et al., 2020; Liu et al., 2021a, b; Luo et al., 2020; Ma et al., 2025; Yang et al., 2025b, 2023; Wu et al., 2016; Wang et al., 2025) have been proposed to leverage tiered memory systems to improve application performance.

### v2

@@ Abstract.
Message Passing Interface (MPI) is a foundational programming model for high-performance computing. MPI libraries traditionally employ network interconnects (e.g., Ethernet and InfiniBand) and network protocols (e.g., TCP and RoCE) with complex software stacks for cross-node communication. This paper presents cMPI 111cMPI is open-sourced at https://github.com/skhynix/cMPI., the first work to optimize MPI point-to-point communication (both one-sided and two-sided) using CXL memory sharing on a real CXL platform, transforming cross-node communication into memory transactions and data copies within CXL memory, bypassing traditional network protocols. We analyze performance across various interconnects and find that CXL memory sharing achieves 7.2-8.1 lower latency than TCP-based interconnects deployed in small- and medium-scale clusters. We address challenges of CXL memory sharing for MPI communication, including data object management over the dax representation (ma2024hydrarpc), cache coherence, and atomic operations. Overall, cMPI outperforms TCP over standard Ethernet NIC and high-end SmartNIC by up to 49 and 72 in latency and bandwidth, respectively, for small messages.
@@ 1. Introduction
Given two MPI processes distributed on two nodes, two types of MPI communication can be employed, either one-sided or two-sided depending on process participation. The MPI communication traditionally employs high-performance interconnect (e.g., Ethernet), and relies on Remote Direct Memory Access (RDMA) or network protocols (e.g., RoCE (10.1145/3543176)). The support of high-performance interconnect includes a complicated software stack, including application-level protocol implementation and system-level driver for device registration and interrupt handling. Using the interconnect technologies also requires the MPI library to use special APIs or programming methods (e.g., socket programming).
@@
Recent work (10.1145/3488423.3519336; 10.1145/3581784.3607108; 10.1007/978-3-031-32041-5_16) reveals low memory utilization in individual nodes in data centers or (supercomputers), and demonstrates the potential of memory disaggregation for improving memory utilization and reducing production cost on supercomputers (10.1145/3624062.3624175; 10.1145/3581784.3607108; 10.1145/3575693.3578835). The emerging Compute Express Link (CXL) (sharma2023introduction) makes memory disaggregation highly feasible. CXL is an interconnect technology designed for high-speed communication among processors, accelerators, and memory devices. CXL allows for memory expansion beyond traditional CPU memory channels by enabling a CPU to communicate with DRAM managed by its own controller over a PCIe channel, hence offering increased capacity and bandwidth without increasing the number of primary CPU memory channels. The CXL memory can be accessed using traditional CPU instructions (e.g., load and store), allowing the existing applications to use it with little changes. The CXL memory has been explored as a memory tier to build HPC systems with larger memory capacity for checkpointing or accommodating scientific applications with larger memory footprint (10024061; 10.1145/3624062.3624175; 10898637; wang2025cxl_performance; 10.1145/3581784.3607108).
@@
In this paper, we study a new MPI communication scheme using CXL memory sharing. The CXL memory sharing is based upon the CXL pooled memory platform (a solution for memory disaggregation), shown in Figure 1. Using the CXL memory sharing, the MPI communication across nodes can be transformed into regular memory transactions and data copy within the CXL memory, without going through the traditional network protocols. There are a couple of benefits of doing so. First, the communication latency can be shorter, compared with network‑based approaches such as TCP over Ethernet and RoCEv2. For example, given a message size of 8 bytes, we show that using a real CXL pooled-memory platform (not based on hardware simulation or emulation (ahn2022enabling; memsys_cxl_switch_2024; osu_omb)) to access the CXL memory takes only 790 , while using TCP over a standard Ethernet NIC and TCP over Mellanox (CX-6 Dx) to access data across nodes take 16 and 18 respectively, about an order of magnitude longer than using the CXL memory sharing (see Section 2 for details). Second, using the CXL memory sharing leads to a simpler software stack, compared to using the traditional network protocols. This is because the CXL protocol to perform CXL memory transactions for data (de)packaging and synchronization between CPU and memory device are done by hardware (not system software).
@@
However, using the CXL memory sharing for MPI communication across nodes faces challenges. First, the representation of the CXL memory sharing in the host (node) creates challenges to manage data objects in the CXL memory. At the node, the CXL memory sharing is exposed as a Direct Access (dax) device by the CXL driver and daxctl (ma2024hydrarpc). The dax device bypasses the traditional page cache and provides memory-level access. However, the abstract of the dax device cannot provide flexibility to create arbitrary data objects and manage their life cycle (e.g., creation and deletion).
@@
We develop a library, cMPI, by extending MPICH-4.2.3 (MPICH_online). We address the limitation of CXL memory sharing to support MPI communications.
@@ 2.2. CXL
The recent CXL specification (CXL 3.0) introduces CXL memory sharing. Based upon the Multi-Headed Device (MHD) and Dynamic Capacity Device (DCD), a host (node) can connect to the CXL pooled memory platform through a dedicated CXL port, ensuring bandwidth fairness and eliminating added latency of a switch hop, shown in Figure 1. We use a CXL pooled memory platform, Niagara 2.0 (Niagara20), for CXL memory sharing.
@@
To measure latency and bandwidth, we use Intel MLC (mlc) for the cases (1) and (7), iPerf (iPerf_online) for (2) and (3), and Perftest (perftest) for (4). We choose those different tools (benchmarks) based on whether they are aligned with the specific system features (e.g., RDMA, or TCP over Ethernet). As we do not have Mellanox CX-3 and CX-6, the performances for the cases (5) and (6) are taken from the vendor’s product report (MellanoxCX3; MellanoxCX6). For the case (8), since existing tools do not support testing memory access with cache flushing on a dax device (the system-level abstraction for CXL memory sharing), we develop a micro-benchmark. This benchmark uses mmap to obtain the virtual address space of the dax device. For latency tests, this benchmark performs memset with cache flushing. During the test, we vary the data size and measure the average latency over 1000 iterations. For bandwidth tests, this benchmark uses multiple threads, and each thread accesses 512 bytes with cache flushing on a dax device region. We measures aggregated bandwidth. See Table 1 for results.
@@ 3.1. CXL SHM Arena
Multi-level hashing. For efficient SHM object indexing, we employ a hash-based structure rather than a plain array to construct the meta-data region. We design the hash structure based on two requirements. First, it should effectively resolve hash collisions while maintaining a fixed capacity to avoid dynamic resizing. Second, it must support parallel insertions and searches. To meet these requirements, we use an existing fixed-capacity multi-level hashing scheme (broder1990multilevel), which is a representative of many multi-level hashing techniques (broder1990multilevel; broder2001using; zuo2019level; xiong2023dalea; zuo2018write; chen2020lock).
@@
Famfs vs. CXL SHM Arena. Famfs (famfs2024) is a shared-memory file system framework recently created by MICRON. Famfs is different from our design from three perspectives. First, Famfs lives in the Linux kernel. In contrast, CXL SHM Arena is a user-space library without any kernel modification. Second, Famfs uses a client/master architecture, and restricts the creation of shared files exclusively to the master node, limiting its applicability to MPI where any node may need to create SHM objects. In contrast, CXL SHM Arena supports the creation of SHM objects by arbitrary nodes. Third, the SHM management APIs in CXL SHM Arena resemble those in the traditional POSIX SHM, which facilitates easy integration into the application, whereas Famfs uses quite different APIs.
@@ 3.5. Memory Coherence
CXL SHM allows multiple hosts to access the same memory region. However, our CXL SHM lacks the cache coherence support, which causes the updates made by one host to be invisible to others. To maintain data coherence, we must rely on either non-cacheable memory sharing or a software-based cache coherence mechanism. While CXL 3.0 introduces support for Back-Invalidate (BI) to enforce hardware coherence across hosts, the anticipated growth in CXL memory size makes impractical a precise snoop filter that tracks each cache line (jain2024memory). In contrast, the software-based cache coherence has been shown to succeed to optimize performance in large shared address space (CXLReady_Micro23; ma2024hydrarpc). We discuss non-cacheable memory sharing and software-based cache coherence as follows.
@@
Software-based cache coherence. We use cache flush and memory fence. Cache flush instructions, such as clwb, clflush, and clflushopt, invalidate the cache line for specific addresses from each level of the cache hierarchy in the coherence domain; if that cache line contains modified data, that data is written back to memory (guide2011intel). Memory fence instructions, such as store fence (sfence) and load fence (lfence), enforce ordering constraints on memory operations, ensuring that all operations issued prior to the fence are completed before those issued after the fence. We perform cache flush and memory fence when accessing the CXL SHM Arena, CXL SHM-based RMA windows, and CXL SHM-based message queues. Specifically, after every write, we perform a cache flush followed by a memory fence to guarantee that data is written back to memory. Conversely, before every read, we perform a memory fence followed by a cache flush to ensure that stale data, potentially held in caches or due to prefetching, is invalidated.
@@ 3.6. Discussion
cMPI currently focuses on one-sided and two-sided communications across nodes. There are collective communications such as Allgather, Allreduce, and ReduceScatter that can benefit from CXL SHM. Within an MPI library, these collective communications are often implemented using point-to-point communications based on certain algorithms (e.g., recursive doubling (arap2015adaptive) and Bruck’s algorithm (fan2022optimizing)), hence the collective communications can directly benefit from cMPI. Nevertheless, there are multiple challenges of using cMPI for the collective communicators, such as communicator handling and integrating intra/inter-node point-to-point communications. We leave collective communications as future work.
@@ 3.7. More Implementation Details
We implement cMPI by extending MPICH-4.2.3 (mpich2). In CXL SHM Arena, to enable cMPI to create millions of distinct CXL SHM objects, we configure a 10-level multi-level hash table and set the maximum number of slots at the level one to 200,000. Since the number of slots is constrained to be a prime number, the actual slot counts across the levels 1-10 range from 199,999 to 199,873. This results in a total of 1,999,260 slots across all levels.
@@ 4.1. Evaluation Methodology
Real CXL Platform. We use Niagara 2.0 (Niagara20), an FPGA-based CXL pooled memory platform. This platform can connect up to 4 nodes without a CXL switch. Each server connects to this platform via an MCIO x8 cable using PCIe 4.0 (16 GT/s per lane). The platform uses four DDR4-2400 DIMM channels, with a total capacity of 1TB and a bandwidth of 19.2 GB/s per channel.
@@
Simulator for scalability study. We use SimGrid (CASANOVA2025103125), an event-based simulator that can study MPI applications with user-specified interconnect latency and bandwidth. Because our CXL platform is constrained to connecting only four hosts, simulation is necessary to study scalability beyond this hardware limitation.
@@ 4.2. Microbenchmarks
We use the OSU Micro-Benchmark suite (OMB) (osu_omb) to compare performance (i.e., latency and bandwidth) across two nodes using CXL SHM, TCP over Ethernet, and TCP over Mellanox (CX-6 Dx). To evaluate the performance of one-sided communication under multiple processes, we extend the OMB’s benchmark for one-sided communication, which originally only supports two processes, to support any number of MPI origin and target processes. To evaluate the performance of two-sided communication, we change the size per message cell to 64KB to improve performance (see Section 4.3). We evaluate the message size from 1B to 8MB, representing small, medium, and large message sizes (osu_omb). For one-sided communication, given MPI processes, half of them is used as the origin processes, and the other half is used as the target processes.
@@ 4.4. Scalability Study
We use CG (a conjugate gradient solver) from NAS Parallel Benchmark (NPB) Suite (doi:10.1177/109434209100500306) 3.4. For CG, we use Class D as input. We also use miniAMR (miniAMR) (a proxy application for adaptive mesh refinement) with the input block size 4 in directions. We choose CG and miniAMR for evaluation, because their communications are dominated by MPI two-sided communications (osti_10450791; CG_p2p).
@@ 5. Related Work
Intra-node MPI communication over shared memory. Optimizing intra-node MPI communication within shared memory has been extensively studied (10.1145/2462902.2462903; 10.1007/978-3-540-87475-1_21; 4663761; 10.1007/978-3-031-73370-3_6; 10898940; memoManaMPI_EuroMPI17; white2020optimizing; adaptiveMPI). For example, Adam et al. (10.1007/978-3-031-73370-3_6) reveal significant bandwidth differences between cores and sockets in a node and demonstrate that current MPI implementations may not fully utilize the intra-node bandwidth potential. Walia et al. (10898940) leverage OpenMP for accelerating MPI_Allreduce in deep learning workloads. Cho et al. (memoManaMPI_EuroMPI17) improve MPI bandwidth performance through on-package memory and huge page support. White and Kale (white2020optimizing) optimize point-to-point communication in AMPI (adaptiveMPI) by their locality-aware design in shared memory. Unlike previous works focusing on intra-node communication, our study optimizes inter-node MPI communication over CXL SHM.
@@
Evaluation of CXL. Recent CXL memory evaluations provide insights into this emerging memory technology (liu2024dissectingcxlmemoryperformance; CXLReady_Micro23; yang2025architecturalimplicationscxlenabledtiered; wang2025cxl_performance; CXLReady_Micro23; PerfCostCXL_Eurosys24; memsys_cxl_switch_2024). For example, Sun et al. (CXLReady_Micro23) demonstrate important differences between real and emulated CXL memory, necessitating reconsideration of previous assumptions on the CXL hardware. Tang et al. (PerfCostCXL_Eurosys24) evaluate CXL performance and design an Abstract Cost Model to estimate the cost-benefit of using CXL memory in datacenters. Wang et al. (wang2025cxl_performance) provide a comprehensive evaluation of how CXL interacts with GPU in LLM training and inference.
@@
CXL shared memory management. Famfs (famfs2024; 10898426) from Micron is an open source file system for scale-out disaggregated shared memory across multiple nodes. We discuss its difference from cMPI in Section 3.1. Zhang et al. (FailureCXL_sosp23) develop a failure-resilient CXL SHM system using reference counting to tolerate partial client failures.
@@
CXL for MPI. OMB-CXL (OMBCXL2024) introduces a micro-benchmark suite for evaluating MPI communications over CXL SHM based on emulation. Ahn et al. (10825804) present the first MPI over CXL SHM implementation beyond benchmarking, utilizing CXL SHM to accelerate MPI_Allgather across nodes. However, this work addresses only a single collective operation and is based on emulation. In contrast, our work is the first to optimize MPI point-to-point communication (both one-sided and two-sided) on real CXL Pooled Memory Platforms, with novel approaches to build DAX abstraction, MPI synchronization, and memory coherence.
@@
Applications of CXL memory. CXL has been explored by various application domains. For databases, Ahn et al. (ahn2022enabling) leverage CXL memory expansion to enhance in-memory database systems. For AI workloads, Xie et al. (SmartQuant) introduce SmartQuant, a CXL-based AI model store supporting runtime weight quantization to reduce inference latency. For approximate nearest neighbor search (ANNS), Jang et al. (CXLANNSATC23) propose CXL-ANNS, which utilizes CXL-based memory disaggregation and parallel search techniques to efficiently handle large datasets.
@@
Tiered memory systems. Apart from serving as a shared memory pool, CXL can also be used to build a tiered memory system for a host. With a tiered memory system, various memory components (wang2020characterizing; xiang2022characterizing; bergman2022reconsidering; dragojevic2014farm; gu2017efficient; abulila2019flatflash; van2019hoti; sharma2022compute; gpu_dp_micro14; wang2025cxl_performance) exhibit differences in latency, bandwidth, capacity, and cost. A range of solutions (Maruf2022MULTICLOCKDT; Kumar2021RadiantEP; Hildebrand2020AutoTMAT; Li2022TransparentAL; Ren2021SentinelET; Wang2019PantheraHM; Dulloor2016DataTI; Raybuck2021HeMemST; Kannan2017HeteroOSO; Kim2021ExploringTD; Weiner2022TMOTM; sc18:wu; unimem:sc17; ics21:warpx; wahlgren2023quantitative; ppopp23:merchandiser; eurosys24:mtm; atc24_hm; micro18:pim; 8048928; neurips20:hm-ann; ics21:athena; ppopp21:sparta; luo:NGS; hpca25:dlrm; hpca25:buffalo; betty:asplos23; hpdc16:wu; wang2025cxl_performance) have been proposed to leverage tiered memory systems to improve application performance.